# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

### v6

@@ Abstract
Learning to execute long-horizon mobile manipulation tasks is crucial for advancing robotics in household and workplace settings. However, current approaches are typically data-inefficient, underscoring the need for improved models that require realistically sized benchmarks to evaluate their efficiency. To address this, we introduce the LAMBDA () benchmark111lambdabenchmark.github.ioââLong-horizon Actions for Mobile-manipulation Benchmarking of Directed Activitiesââwhich evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection. Our benchmark includes human-collected demonstrations that provide realism and diversity in simulated and real-world settings. Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation. We leverage LAMBDA to benchmark current end-to-end learning methods and a modular neuro-symbolic approaches that combines foundation models with task and motion planning. We find that end-to-end methodsâeven when pretrainedâyield lower success rates, while neuro-symbolic methods perform significantly better and require less data.
@@ I INTRODUCTION
Improving data efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications. These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions. Significant strides have been made in tabletop manipulation with various approaches for data-efficient learning [1, 2, 3], however existing MoMa models typically rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6]. Collecting such large-scale datasets is resource-intensive, requiring substantial time, labor, and infrastructure. This data inefficiency arises from the inherent challenges of MoMa tasks: large observation and state spaces, diverse tasks, and the need for generalization across environments.
@@
Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, leaving a significant gap in evaluating progress in this area. A few works  [7, 8, 9, 10, 4, 11] pair task demonstrations with language instruction annotations and support long-horizon tasks [7, 12, 8, 11], providing an intuitive and accessible interface for specifying robot objectives. Even fewer [9, 11] provide natural language instructions in an open-ended, flexible format rather than relying on fixed templates. Many benchmarks lack real-world data [13, 12, 9, 14, 15, 4] or rely on planner-generated trajectories [16, 10, 17], limiting their practicality. Quadrupeds excel in diverse terrains [18] and are well-suited for long-horizon MoMa tasks, yet only RoboSuite [19], RoboCasa [9], and DeMoBot [15] provide quadruped data.
@@
To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities). evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks. Our dataset consists of trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability. Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments. By including language-conditioned tasks, supports intuitive and accessible interactions, making it practical for real-world users. Unlike planner-generated data, human-collected trajectories in provide natural variability and expertise, better capturing real-world robot operations. Additionally, supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios. Simulated environments enable diverse testing scenarios and object variability, while real-world data ensures models are prepared for practical deployment. We collected this dataset with a quadruped robot [18] due to its optimal form-factor suitable for navigating complex terrain.
@@
We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the datasetâs ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems.
@@
Overall our contributions are summarized as follows: (1) benchmark:We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings. (2) High-quality dataset:âs tasks are paired with human-collected robot demonstrations, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation. (3) Comprehensive evaluation:We benchmark several models, including BC and modular neuro-symbolic approaches, providing a baseline for future model development for MoMa tasks.
@@ II Related Work
As robot learning algorithms evolve rapidly, benchmarks have become critical for evaluating algorithmic decisions and ensuring consistent performance comparisons in standardized scenarios. However, existing benchmarks, shown in Table I, fail to adequately evaluate models for data efficiency in long-horizon, language-conditioned indoor MoMa tasks, leaving significant gaps in advancing this area. For example, most benchmarks [20, 13, 12, 16, 19, 14, 15, 21, 22], lack natural language annotations, despite its importance as an intuitive interface for specifying tasks, particularly in household and workplace scenarios. Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications. Even fewer benchmarks, such as RoboCasa [9] and PARTNR [11], offer natural language instructions in a free-form, unconstrained format rather than predefined templates. This flexibility is crucial, as it better reflects the variability and nuance of human communication that robots will encounter in-the-wild.
@@
Moreover, almost all benchmarks [20, 13, 16, 19, 9, 14, 15, 21, 10, 22, 4, 17], do not support long-horizon tasks involving navigation across rooms and floors. These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations. Without this capability, benchmarks fail to reflect the complexity of real-world scenarios. As shown in Table I, Mobile ALOHA [13] and BRMData [14] lack simulation data entirely. Simulated environments offer diverse scenes and objects, making it easier to test models across a wide range of scenarios. Their standardized and controlled conditions create a portable testbed that others can readily use, ensuring consistency and reproducibility in evaluations. On the other hand, most lack real-world data [7, 20, 16, 8, 19, 21, 10, 22, 11, 17], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
@@
The majority of benchmarks in Table I provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models. However, a few, SafeAgentBench [7], HumanoidBench [20], BEHAVIOR-1K [12], InfiniteWorld [8], and PARTNR [11] lack this attribute, limiting their utility for model training and evaluation. M3Bench [10], M2Diffuser [17], and ManiSkill-HAB [16] depend on planner-generated trajectories. These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
@@
Finally, as shown in Table I, RoboCasa [9], RoboSuite [19], and DeMoBot [15] are the only benchmarks featuring armed quadruped data, which is essential for tasks requiring stability and adaptability. Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as the ones present in . Their robustness as quadrupeds allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios. The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
@@
By addressing these limitations, the benchmark provides a comprehensive benchmark to test the data efficiency of MoMa models. It integrates language-conditioned tasks, long-horizon trajectories spanning multiple rooms and floors, and both simulated and real-world data. Additionally, its human-collected demonstrations ensure realistic variability, and its inclusion of armed quadruped data makes it uniquely suited for advancing robotics in complex real-world settings.
@@ III Benchmark
The LAMBDA () benchmark is designed to evaluate model data efficiency on long-horizon MoMa tasks. features language-conditioned, multi-room, multi-floor pick-and-place tasks that span a wide range of diverse environments across both simulated and real settings. The tasks include their respective human-collected demonstrations, capturing natural variability and providing realistic scenarios for training and evaluation.
@@ III-A Challenges of Tasks
Data Efficiency in MoMa. We define data efficiency as the ability to achieve relatively strong performance with minimal robot trajectory demonstrations, where each trajectory consists of a sequence of robot observations, states, and actions at every time step. Non-end-to-end systems are still considered within our definition of data efficiency. For example, modular methods that integrate off-the-shelf Vision-Language Models (VLMs), which are not trained on robot demonstration trajectories, are not excluded on this basis. Unlike very large MoMa datasets such as RT-1 [4], is designed to test the data efficiency of learning-based models, done in Section IV, by offering a dataset of manageable size, as shown in Table II.
@@
While previous tabletop manipulation methods depended on thousands of demonstrations [23], recent models [1, 2, 3] are more data efficient, requiring only tens of demonstrations. In a similar direction, recognizing the increased complexity of MoMa, we provide hundreds of demonstrations for model development and benchmarking. This deliberate size balances diversity for learning with a manageable scale while also challenging models to be more data efficient, avoiding the excessive resource costs of unrealistically large datasets.
@@
Long Horizon Tasks in Complex Spaces. MoMa tasks introduce significant complexity by involving larger and more intricate action, observation, state, and task spaces compared to tabletop manipulation. Operating in dynamic three-dimensional environments that span multiple rooms and floors (shown in Figure LABEL:fig:main_splash), these tasks require the seamless integration of high-level planning with low-level motion control. For example, an agent may need to navigate to a dining table, pick up a plate, and transport it to a countertop in another room. Even in this seemingly straightforward pick-and-place scenario, the nature of the task manifests in various approachesâthe agent might approach the table from different angles, select from multiple grasping points on the plate [24], or choose among several routes to the destination.
@@
The challenges are further amplified in long-horizon tasks, where extended action sequences necessitate hierarchical strategies [25] that consider spatial relationships, sequential dependencies, and the compounding effects of errors over timeâissues that are also observed in covariate shifts with behavioral cloning approaches [26]. Unlike fixed single-arm setups, MoMaâs expansive spaces force the agent to contend with a higher degree of variability and uncertainty. Consequently, our benchmark focuses on pick-and-place tasks, such as the ones presented in Figure 1, which serve as a representative foundation for addressing the broader demands of MoMa models in complex, real-world environments.
@@
Crowdsourced Free-Form Natural Language Instructions. Understanding language is crucial for robots in human environments, as it is the most natural way for humans to communicate task specifications. Thus, âs tasks are language-conditioned, meaning goals are not specified via images, state information, or symbolic representations. Language adds complexity, requiring models to interpret and ground linguistic expressions [27] instead of relying on explicit visual or spatial cues.
@@
To ensure tasks in reflect real-world needs, we collected free-form human-generated instructions via crowdsourcing. This provides a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings. Unlike system-generated templated commands, âs free-form instructions, seen in Figures LABEL:fig:main_splash & 1, vary in phrasing, detail, and structure, requiring models to generalize across a variety of expressions rather than memorizing fixed patterns. This pushes models to interpret and reason about task objectives. Additionally, these instructions introduce ambiguities and implicit references further testing a modelâs ability to align interpretations with human intent in long-horizon MoMa tasks.
@@ III-B Simulated Dataset
The dataset for our simulated tasks consists of trajectories, with an average length of timesteps, collected across rooms in five distinct environments, designed to ensure diversity in both object types and room layouts. We utilized the AI2THORÂ [28] simulator, specifically the RoboTHORÂ [29] environments, which feature connected rooms suitable for long-horizon tasks.
@@
To generate realistic and diverse instructions, we leveraged Prolific222prolific.com for crowdsourcing to gather natural language commands from participants, reflecting tasks that people would naturally expect robots to perform in indoor settings. We teleoperated the simulated robot with a custom keyboard-based teleoperation system we developed to execute these commands. After teleoperation, each trajectory was replay-verified to confirm correctness, ensuring data integrity.
@@ III-C Real-World Dataset
The dataset for our real-world tasks comprises trajectories collected across rooms in three distinct environments. These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities. The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building. These include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to challenge model generalizability.
To handle these complex terrains, we teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on natural language commands provided by seven participants. The commands, distributed as , , and across the three environments, respectively, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings. The trajectories were collected at Hz.

### v7

@@ Abstract
Learning to execute long-horizon mobile manipulation tasks is crucial for advancing robotics in household and workplace settings. However, current approaches are typically data-inefficient, underscoring the need for improved models that require realistically sized benchmarks to evaluate their efficiency. To address this, we introduce the LAMBDA () benchmark111lambdabenchmark.github.ioââLong-horizon Actions for Mobile-manipulation Benchmarking of Directed Activitiesââwhich evaluates the data efficiency of models on language-conditioned, long-horizon, multi-room, multi-floor, pick-and-place tasks using a dataset of manageable size, more feasible for collection. Our benchmark includes human-collected demonstrations that provide realism and diversity in simulated and real-world settings. Unlike planner-generated data, these trajectories offer natural variability and replay-verifiability, ensuring robust learning and evaluation. We leverage to benchmark current end-to-end learning methods and a modular neuro-symbolic approach that combines foundation models with task and motion planning. We find that learning methods, even when pretrained, yield lower success rates, while a neuro-symbolic method performs significantly better and requires less data.
@@ I INTRODUCTION
Improving data-efficient learning for long-horizon mobile manipulation (MoMa) tasks is vital for pushing the boundaries of robotics in household and workplace applications. These tasks, such as fetching an item from a specific location or organizing objects across multiple rooms, often involve understanding language commands, navigating complex spaces, and performing precise pick-and-place actions. Strides have been made in tabletop manipulation with various approaches for data-efficient learning [1, 2, 3]; however, existing MoMa models typically rely on large-scale datasets with extensive amounts of trajectories [4, 5, 6]. Collecting such large-scale datasets is resource-intensive, requiring substantial time, labor, and infrastructure. This data inefficiency arises from the inherent challenges of MoMa tasks: extensive observation, action, and state spaces, as well as diverse tasks, and the need for generalization across environments.
@@
Existing benchmarks do not assess models for data efficiency, as they rely on unrealistically large sizes, leaving a significant gap in evaluating progress in this area. A few works  [7, 8, 9, 10, 4, 11] pair task demonstrations with language instruction annotations and support long-horizon tasks [7, 12, 8, 11], providing an intuitive and accessible interface for specifying robot objectives. Even fewer [9, 11] provide natural language instructions in an open-ended, flexible format rather than relying on fixed templates. Many benchmarks lack real-world data [13, 12, 9, 14, 15, 4] or rely on planner-generated trajectories [16, 10, 17], limiting their practicality. Quadrupeds excel in diverse terrains [18] and are well-suited for long-horizon MoMa tasks, yet only RoboSuite [19], RoboCasa [9], and DeMoBot [15] provide real quadruped data.
@@
To address these challenges, we propose the LAMBDA () benchmark (Long-horizon Actions for Mobile-manipulation Benchmarking of Directed Activities). evaluates models on a realistic amount of demonstrations of language-conditioned, multi-room, multi-floor, pick-and-place tasks requiring long-horizon planning and execution, addressing the limitations of existing benchmarks. Our dataset consists of trajectories, which were crowdsourced and manually collected to ensure realism, diversity, and replay-verifiability. Crowdsourcing was specifically conducted to capture natural, flowing instructions rather than templated commands, providing a realistic distribution of tasks that people want robots to perform in household and workplace environments. By including language-conditioned tasks, supports intuitive and accessible interactions, making it practical for real-world users. Unlike planner-generated data, human-collected trajectories in provide natural variability and expertise, better capturing real-world robot operations. Additionally, supports long-horizon tasks involving navigation across rooms and floors, reflecting the realistic demands of indoor scenarios. Simulated environments enable diverse testing scenarios and object variability, while real-world data ensures models are prepared for practical deployment. We collected the real data with a quadruped robot due to its optimal form factor, suitable for navigating complex terrain.
@@
We used to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, such as and , indicating data inefficiency while confirming the datasetâs ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved a success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems.
@@
Overall, our contributions are summarized as follows: (1) benchmark:We introduce a novel benchmark specifically tailored in size to evaluate data efficiency in MoMa, emphasizing language-conditioned, long-horizon pick-and-place indoor tasks in realistic household and workplace settings. (2) High-quality dataset:âs tasks are paired with human-collected robot demonstrations, providing natural variability, replay-verifiability, and a balance of simulated and real-world data for robust model evaluation. (3) Comprehensive evaluation:We benchmark multiple methods, including BC and modular neuro-symbolic approaches, providing a baseline for future MoMa system development.
@@ II RELATED WORK
As robot learning algorithms evolve rapidly, benchmarks have become critical for evaluating algorithmic decisions and ensuring consistent performance comparisons in standardized scenarios. However, existing benchmarks, shown in Table I, fail to adequately evaluate models for data efficiency in long-horizon, language-conditioned indoor MoMa tasks, leaving significant gaps in advancing this area. For example, most benchmarks [20, 13, 12, 16, 19, 14, 15, 21, 22], lack natural language annotations, despite its importance as an intuitive interface for specifying tasks, particularly in household and workplace scenarios. Language conditioning allows robots to interpret human commands directly, making this an essential feature for realistic applications. Even fewer benchmarks, such as RoboCasa [9] and PARTNR [11], offer natural language instructions in a free-form, unconstrained format rather than predefined templates. This flexibility is crucial, as it better reflects the variability and nuance of human communication that robots will encounter in the wild.
@@
Moreover, almost all benchmarks [20, 13, 16, 19, 9, 14, 15, 21, 10, 22, 4, 17] do not support long-horizon tasks involving navigation across rooms and/or floors. These tasks are critical for indoor robotics, where robots must traverse long distances to perform daily operations. Without this capability, benchmarks fail to reflect the complexity of real-world scenarios. As shown in Table I, Mobile ALOHA [13] and BRMData [14] lack simulation data entirely. Simulated environments offer diverse scenes and objects, making it easier to test models across a wide range of scenarios. Their standardized and controlled conditions create a portable testbed that others can readily use, ensuring consistency and reproducibility in evaluations. On the other hand, most lack real-world data [7, 20, 16, 8, 19, 21, 10, 22, 11, 17], which is essential for providing a benchmark grounded in realistic scenarios where robots will ultimately be deployed, ensuring models are tested under practical and uncontrolled conditions.
@@
The majority of benchmarks in Table I provide expert demonstrations for the tasks, which are crucial for establishing ground-truth data to train learning-based methods and test models. However, a few, SafeAgentBench [7], HumanoidBench [20], BEHAVIOR-1K [12], InfiniteWorld [8], and PARTNR [11], lack this attribute, limiting their utility for model training and evaluation. M3Bench [10], M2Diffuser [17], and ManiSkill-HAB [16] depend on planner-generated trajectories. These lack the natural variability and expertise of human-collected demonstrations, which more accurately reflect real-world conditions and improve the quality of learned behaviors by providing nuanced, high-quality examples that enable better generalization.
@@
Finally, as shown in Table I, RoboCasa [9], RoboSuite [19], and DeMoBot [15] are the only benchmarks featuring armed real quadruped data, which is essential for tasks requiring stability and adaptability. Quadrupeds with manipulators can handle uneven terrains, such as stairs, while carrying objects, making them ideal for multi-floor environments, such as those found in . Their robustness allows them to maintain stability in obstacle-dense environments, withstand collisions, and recover by picking themselves up if they fall, making them especially suitable for realistic and demanding scenarios. The absence of this embodiment limits the applicability of the vast majority of benchmarks to realistic robotic use cases.
@@
By addressing these limitations, provides a comprehensive benchmark to test the data efficiency of MoMa models. It integrates language-conditioned tasks, long-horizon trajectories spanning multiple rooms and floors, and both simulated and real-world data. Additionally, its human-collected demonstrations ensure realistic variability, and its inclusion of armed real quadruped data makes it uniquely suited for advancing robotics in complex settings.
@@ III BENCHMARK
The LAMBDA () benchmark is designed to evaluate model data efficiency on long-horizon MoMa tasks. features language-conditioned, multi-room, multi-floor pick-and-place demonstrations spanning a range of diverse environments across simulated and real settings.
@@ III-A Challenges of Tasks
Data Efficiency in MoMa. We define data efficiency as the ability to achieve relatively strong performance with minimal robot demonstrations, where each trajectory consists of a sequence of robot observations, states, and actions at every timestep. Non-end-to-end systems are still considered within our definition of data efficiency. For example, modular methods that integrate off-the-shelf Vision-Language Models (VLMs), which are not trained on robot demonstrations, are not excluded on this basis. Unlike large MoMa datasets such as RT-1 [4], is designed to test the data efficiency of learning-based models, done in Section IV, by offering a dataset of manageable size, as shown in Table II.
@@
While previous tabletop manipulation methods depended on thousands of demonstrations [23], recent models [1, 2, 3] are more data efficient, requiring only tens of demonstrations. In a similar direction, recognizing the increased complexity of MoMa, we provide hundreds of demonstrations for model development and benchmarking. This deliberate size balances diversity for learning with a manageable scale while also challenging models to be more data efficient in out-of-distribution generalization, avoiding the excessive resource costs of unrealistically large datasets.
@@
Long Horizon Tasks in Complex Spaces. MoMa tasks introduce significant complexity by involving larger and more intricate action, observation, state, and task spaces compared to tabletop manipulation. Operating in environments that span multiple rooms and floors (Figure LABEL:fig:main_splash), these tasks require the integration of high-level planning with low-level motion control. For example, a robot may need to navigate to a dining table, pick up a plate, and transport it to a countertop in another room. Even in this seemingly straightforward pick-and-place scenario, the nature of the task manifests in various approachesâthe robot might approach the table from different angles, select from multiple grasping points on the plate [24], or choose among several routes to the destination.
@@
The challenges are further amplified in long-horizon tasks, where extended action sequences necessitate hierarchical strategies [25] that consider spatial relationships, temporal dependencies, and the compounding effects of errors over timeâissues that are also observed in covariate shifts with BC approaches [26]. Unlike fixed single-arm setups, MoMaâs expansive spaces force the robot to contend with a higher degree of variability and uncertainty. Consequently, our benchmark focuses on pick-and-place tasks (e.g., Figure 1), which serve as a representative foundation for addressing the broader demands of MoMa models in complex environments.
@@
Crowdsourced Free-Form Natural Language Instructions. Understanding language is crucial for robots in human environments, as it is the most natural way for humans to specify tasks. Thus, âs tasks are language-conditioned, meaning goals are not specified via images, state information, or symbolic representations. Language adds complexity, requiring models to interpret and ground linguistic expressions [27] instead of relying on explicit visual or spatial cues.
@@
To ensure tasks in reflect real-world needs, we collected free-form human instructions via crowdsourcing. This provides a diverse and realistic set of tasks that people expect robots to perform in household and workplace settings. Unlike system-generated templated commands, âs free-form instructions (e.g., Figures LABEL:fig:main_splash & 1) vary in phrasing, detail, and structure, requiring models to generalize across variations rather than memorizing fixed patterns. This pushes models to interpret and reason about task objectives. Additionally, these instructions introduce ambiguities and implicit references, further testing a modelâs ability to align interpretations with human intent in long-horizon MoMa tasks.
@@ III-B Simulated Dataset
The dataset for our simulated tasks consists of trajectories, with an average length of timesteps, collected across rooms in five distinct environments, designed to ensure diversity in both object types and unique room layouts. The object configurations within each environment are static. Objects and furniture include, but are not limited to: clocks, bats, basketballs, apples, bowls, plates, forks, couches, TVs, chairs, desks, tables, trash cans, beds, shelves, drawers, lamps, etc. We utilized the AI2THORÂ [28] simulator, specifically the RoboTHORÂ [29] environments, which feature connected rooms suitable for long-horizon tasks.
@@
To obtain realistic and diverse instructions, we used Prolific222prolific.com for crowdsourcing to gather commands from participants, reflecting tasks that people would naturally expect robots to perform in indoor settings. We teleoperated the robot with a keyboard-based teleoperation system we developed, to execute these commands. Afterward, each trajectory was replay-verified to confirm correctness.
@@ III-C Real-World Dataset
The dataset for our real-world tasks comprises trajectories collected across unique rooms in three distinct environments. These environments were chosen for their large size, thereby allowing long-horizon tasks, and their status as workplaces where robots can be particularly useful. The first is a three-room laboratory, the second is a floor in a university building consisting of five rooms, and the third spans two adjacent floors connected by stairs within the same building, with eight rooms. These include kitchens, lobbies, a classroom, and a staircase, offering diverse objects and unique room layouts to challenge model generalizability.
To handle these complex terrains, we teleoperated the Spot [18] robot using a tablet equipped with thumbsticks to collect task trajectories based on natural language commands provided by seven participants. The commands, distributed as , , and across the three environments, respectively, specified room-to-room and floor-to-floor tasks. The trajectories were collected at Hz.